OpenStudy (anonymous):

Ben bought a digital camera for $90. The tax rate is 6%. Which proportion can be used to find the amount of tax Ben will pay? A. Working with Percent B. Working with Percent C. Working with Percent D. Working with Percent

OpenStudy (anonymous):

90 * 6% = 90 * 0,06 = 90 + (90 * 0,,06) =

OpenStudy (anonymous):

the problem is in the attachment

OpenStudy (anonymous):

A

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

can you help me with 1 more?